# When Features Beat Noise: A Feature Selection Technique Through Noise-Based Hypothesis Testing

### v1

@@ 3.1 Population Assumptions
$H_{0j}:Y\;\perp\!\!\!\perp\;f_{j}\quad\text{versus}\quad H_{1j}:Y\;\not\!\perp\!\!\!\perp\;f_{j},$
@@
$H_{0j}:\mathcal{I}(f_{j},Y)=0,\qquad H_{1j}:\mathcal{I}(f_{j},Y)>0.$
@@ 3.2 Approximate solution
To approximate the distribution of importances, we employ the bootstrap. Specifically, we generate $b$ bootstrap replicates of the augmented dataset, each of size $n$. For each replicate, we fit a predictive model—treated as a hyperparameter of the framework—and extract feature importance scores. Denote by $I^{(i)}_{j}(l)$ the importance of candidate feature $f_{j}$ in the $i$-th bootstrap when $l$ noise features are present, and by
@@ 3.3 Algorithmic Summary
Algorithm 1 outlines the proposed Noise-Augmented Bootstrap Feature Selection (NABFS) method.
@@
  1. 1.
    Augment the dataset with $l$ synthetic noise features sampled i.i.d. from $\mathcal{D}_{\varepsilon}$.
  2. 2.
    For each bootstrap replicate $i=1,\dots,b$:
    1. (a)
      Draw a bootstrap sample of size $n$ from $(X,Y)$.
    2. (b)
      Fit the chosen model $\mathcal{M}$ and compute feature importances.
    3. (c)
      Record $I^{(i)}_{j}(l)$ for each candidate feature $f_{j}$ and the maximum noise importance
@@
    4. (d)
      Form the paired difference $D_{i}(l)=I^{(i)}_{j}(l)-M^{(i)}(l)$ for each $f_{j}$.
  3. 3.
    For each feature $f_{j}$, compute the one-sided Wilcoxon signed-rank statistic $T^{+}(l)$ from $\{D_{i}(l)\}_{i=1}^{b}$ and obtain a p-value.
  4. 4.
    Apply the Holm– Bonferroni procedure at level $\alpha$ to adjust p-values.
  5. 5.
    Return $\widehat{\mathcal{S}}=\{f_{j}:\text{adjusted p-value}<\alpha\}$.
Algorithm 1 Noise-Augmented Bootstrap Feature Selection (NABFS)
@@ Appendix A Proofs of Theoretical Results
$p\coloneqq p(l)=|P(l)|,\qquad c_{t}\coloneqq\sum_{j\in N(l)}\mathbf{1}\{|D_{t}(l)|>|D_{j}(l)|\}.$

### v2

# When Features Beat Noise: A Noise-Driven Hypothesis Testing Method for Feature Selection

@@ Abstract
Keywords: Feature selection; Multiple hypothesis testing; Bootstrap inference; Cross-validation; Boruta; Knockoff; Recursive feature elimination; Extra Trees, Hyperparameter tuning.
@@ 3.1 Population Assumptions
$H_{0j}:Y\;\perp\!\!\!\perp\;f_{j}\qquad\text{versus}\qquad H_{1j}:Y\;\not\!\perp\!\!\!\perp\;f_{j},$
@@
$H_{0j}:\mathcal{I}(f_{j},Y)=0,\hskip 18.49988ptH_{1j}:\mathcal{I}(f_{j},Y)>0.$
@@ 3.2 Approximate solution
To approximate the distribution of importances, we employ the bootstrap. Specifically, we generate $b$ bootstrap replicates of the augmented dataset, each of size $n$. For each replicate, we fit a predictive model, treated as a hyperparameter of the framework, and extract feature importance scores. Denote by $I^{(i)}_{j}(l)$ the importance of candidate feature $f_{j}$ in the $i$-th bootstrap when $l$ noise features are present, and by
@@ 3.3 Algorithmic Summary
The algorithm below outlines the proposed Noise-Augmented Bootstrap Feature Selection (NABFS) method.
@@
1. Augment the dataset with $l$ synthetic noise features sampled i.i.d. from $\mathcal{D}_{\varepsilon}$.
2. For each bootstrap replicate $i=1,\dots,b$:
(a) Draw a bootstrap sample of size $n$ from $(X,Y)$.
(b) Fit the chosen model $\mathcal{M}$ and compute feature importances.
(c) Record $I^{(i)}_{j}(l)$ for each candidate feature $f_{j}$ and the maximum noise importance
@@ 2. For each bootstrap replicate $i=1,\dots,b$:
(d) Form the paired difference $D_{i}(l)=I^{(i)}_{j}(l)-M^{(i)}(l)$ for each $f_{j}$.
3. For each feature $f_{j}$, compute the one-sided Wilcoxon signed-rank statistic $T^{+}(l)$ from $\{D_{i}(l)\}_{i=1}^{b}$ and obtain a p-value.
4. Apply the Holm–Bonferroni procedure at level $\alpha$ to adjust p-values.
5. Return $\widehat{\mathcal{S}}=\{f_{j}:\text{adjusted p-value}<\alpha\}$.
@@ Appendix A Proofs of Theoretical Results
$p\coloneqq p(l)=|P(l)|,\hskip 18.49988ptc_{t}\coloneqq\sum_{j\in N(l)}\mathbf{1}\{|D_{t}(l)|>|D_{j}(l)|\}.$